OpenStudy (anonymous):

Exact value of cos75 using identities

OpenStudy (anonymous):

0,2588

OpenStudy (solomonzelman):

cos(75)=cos(45+30) = cos 45 cos 30 - sin 45 sin 30 sin(30)=1/2 cos(30)= sqrt3 /2 sin(45)=cos(45)=1/sqrt2

OpenStudy (solomonzelman):

just finish it of please.

OpenStudy (solomonzelman):

plug in the values I gave you for cos and sin of 30 and 45, and simplify the problem as much as you can. You can type/draw the steps you do and I'll help you.

OpenStudy (anonymous):

I don't understand what you're asking me to do. Attached is an example problem of what I am supposed to do.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

so you needed to find cos(75) using a half-angle formula ?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

can you remind me the half-angle formula for cos, I don;t remember it.

OpenStudy (anonymous):

\[\cos \frac{ u }{ 2 } = \pm \sqrt \frac{ 1 + \cos u }{ 2 }\]

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

connection snapped viewing through profile, anyway....

OpenStudy (solomonzelman):

\[\huge\color{blue}{ Cos ( \frac{150}{2} )=±\sqrt{ \frac{1+\cos(150)}{2} } }\]

OpenStudy (anonymous):

\[\frac{ \sqrt{6} - \sqrt{2} }{ 4 }\] I think this is the correct answer

OpenStudy (solomonzelman):

ok, cos(150)= -sqrt3/2 (trig. table) \[\huge\color{blue}{ Cos ( \frac{150}{2} )=±\sqrt{ \frac{1-\frac{ \sqrt{3}}{2} }{2} } }\] \[\huge\color{blue}{ Cos ( \frac{150}{2} )=±\sqrt{ \frac{ \frac{2}{2}+\frac{ \sqrt{3}}{2} }{2} } }\] sp far, hold...

OpenStudy (solomonzelman):

- , not + in last blue, sorry.

OpenStudy (solomonzelman):